# Plasticine: Accelerating Research in Plasticity-Motivated Deep Reinforcement Learning

### v1

@@ Abstract
Developing lifelong learning agents is crucial for artificial general intelligence. However, deep reinforcement learning (RL) systems often suffer from plasticity loss, where neural networks gradually lose their ability to adapt during training. Despite its significance, this field lacks unified benchmarks and evaluation protocols. We introduce Plasticine, the first open-source framework for benchmarking plasticity optimization in deep RL. Plasticine provides single-file implementations of over 13 mitigation methods, 10 evaluation metrics, and learning scenarios with increasing non-stationarity levels from standard to open-ended environments. This framework enables researchers to systematically quantify plasticity loss, evaluate mitigation strategies, and analyze plasticity dynamics across different contexts. Our documentation, examples, and source code are available at https://github.com/RLE-Foundation/Plasticine.
Keywords: Deep RL, Plasticity, Lifelong Learning, Single-file Implementation
@@ 1 Introduction
Developing agents that never stop learning and evolving in dynamic real-world environments represents a critical frontier for advancing artificial intelligence and a fundamental prerequisite for truly developable systems (Abel et al., 2023). Reinforcement learning (RL), with its inherent properties of learning from experience, continual online improvement, and iterative strategic exploration, emerges as the ideal framework for pursuing this objective (Silver and Sutton, 2025). Despite its promise, recent research has increasingly revealed a critical limitation: plasticity loss, a phenomenon in which neural networks in RL agents gradually lose their ability to adapt and incorporate new information as training progresses (Dohare et al., 2024; Klein et al., 2024), thus significantly impeding the development of truly lifelong learning agents (Lyle and Pascanu, 2024).
@@
Despite the clear importance of mitigating plasticity loss in deep RL, this field lacks unified methodological frameworks and evaluation protocols. This lack of established benchmarks significantly hampers fair comparisons, methodological consistency, and precise problem formulation. Numerous fundamental questions remain inadequately addressed in current research: $\bullet$ How can we reliably measure and quantify plasticity loss through empirical means? $\bullet$ To what degree do existing mitigation approaches effectively resolve this issue? $\bullet$ In what ways does plasticity loss vary across different task domains and learning scenarios? In response to these challenges, we introduce Plasticine, the first open-source framework for benchmarking plasticity research in deep RL, which currently provides high-quality single-file implementations for 13+ methods, 10+ evaluation metrics, and 3+ learning scenarios.
@@ 2 Architecture of Plasticine
Figure 1 illustrates the overall architecture of the Plasticine framework, which contains four core parts: Methods, Metrics, Environments, and Benchmark. To maximize the convenience for researchers, we adopt the “single-file” implementation style that has been widely recognized by the RL community (Huang et al., 2022; Tarasov et al., 2023). More specifically, we put all implementation details of a plasticity loss mitigation method into a single file, including the method logic, hyperparameter configuration, backbone RL algorithm, training environment, and model evaluation. This ensures no intervention exists between different code files and allows researchers to focus on a specific method without worrying about handling special cases while minimizing the workload when developing new methods. In the following content, we introduce the detailed design insights.
Figure 1: Overview of the Plasticine framework. It consists of 4 core parts: The Methods part provides high-quality and single-file implementations of diverse plasticity loss mitigation methods. The Metrics part provides comprehensive evaluation metrics to reflect the plasticity variation during training. The Environments part provides environmental configurations for different levels of continual RL scenarios, where * denotes a variant designed for the continual scenario. Finally, the Benchmark part provides reusable datasets and models for facilitating corresponding research, which is constructed based on Weights & Biases.
@@ 2.1 Plasticity Loss Mitigation Methods
    Reset-based Interventions: Re-initializing portions of the network represents the most direct approach to regain plasticity (Nikishin et al., 2022), whether through manually selecting specific layers or automatically identifying neurons experiencing severe performance plateaus (Sokar et al., 2023; Galashov et al., 2024).
@@
    Optimizer: Specialized optimization approaches like SAM (Lee et al., 2023) and Fast TRAC (Muppidi et al., 2024) help navigate complex loss landscapes during non-stationary training, thereby maintaining plasticity.
@@ 2.2 Comprehensive Evaluation Metrics
2.3 Diverse Evaluation Scenarios: Towards Real Lifelong RL
@@
2.3.1 Level 1: Standard Online RL
Standard online RL naturally exhibits non-stationarity as policies evolve and through bootstrapping processes. In these environments, agents must maintain plasticity throughout training to achieve optimal performance instead of stuck at suboptimal levels. This setting offers a foundation for measuring plasticity during natural distribution shifts in policy optimization without adding explicit task changes. Our baseline evaluation includes popular benchmarks (ALE, Procgen, DMC) that span diverse scenarios with both state-based and vision-based tasks across discrete and continuous action spaces.
2.3.2 Level 2: Continual RL
Continual RL scenarios create more challenging plasticity demands by introducing explicit distribution shifts beyond those naturally occurring in standard RL. We implement two types of scenarios: one where visual features or dynamics models change within the same task, and another requiring adaptation between entirely different tasks. These environments test an agent’s ability to maintain plasticity across both sudden task switches and gradual dynamics changes, positioning them between standard and open-ended learning contexts. With clearly defined non-stationarity boundaries, we can directly observe how models behave before and after distribution shifts, gaining valuable insights into plasticity degradation patterns across different task distributions.
2.3.3 Level 3: Open-Ended RL
Open-ended RL presents the highest level of non-stationarity challenge, closely matching real-world lifelong learning needs. We use Craftax (Matthews et al., 2024) as our evaluation environment, which features complex procedural generation without clear task boundaries. Unlike our previous scenarios with structured transitions, Craftax offers an expansive action space with long-horizon dependencies, hierarchical task structures, and continuous novelty. These environments test the frontier of plasticity maintenance, measuring whether agents can preserve learning abilities while navigating constantly evolving challenges, a key requirement for truly lifelong learning systems.
Appendix A Supplementary Materials
@@
  • Appendix D introduces the included backbone RL algorithms;
  • Appendix E introduces the implementation details of the training environments for the three-level learning scenarios;
  • Appendix F provides formally mathematical definitions of the plasticity-related metrics;
@@ Appendix B Project Structure
[…]
@@ C.5.1 TRAC
Appendix D Base RL Algorithms
In this paper, we select three representative RL baselines as the base algorithms, including policy-based, value-based, on-policy, and off-policy approaches.
D.1 Proximal Policy Optimization
Proximal policy optimization (PPO) (Schulman et al., 2017) is a popular on-policy algorithm for learning a continuous or discrete control policy—$\pi_{\bm{\theta}}(\bm{a}|\bm{s})$. PPO employs policy gradients using action-advantages: $A_{t}=A^{\pi}(\bm{a}_{t},\bm{s}_{t})=Q^{\pi}(\bm{a}_{t},\bm{s}_{t})-V^{\pi}(\bm{s}_{t})$, and minimizes a clipped-ratio loss over mini-batches of recent experience (collected under $\pi_{\bm{\theta}_{\text{old}}}$):
Our PPO agents learn a state-value estimator—$V_{\bm{\phi}}(\bm{s})$, which is regressed against a target of discounted returns and used with generalized advantage estimation (Schulman et al., 2015):
D.2 Parallel Q-Network
Parallel Q-Network (PQN) (Gallici et al., 2025) exploits parallelised sampling without a replay buffer or target networks, which is regularised with network normalization and $\ell_{2}$ regularization. To achieve $\ell_{2}$ regularization, it regularizes the final weight layer $w$ through TD update vector:
$\delta^{\mathrm{reg}}(\phi,\varsigma):=\delta^{\mathrm{Layer}}(\phi,\varsigma)-\left(\frac{\gamma}{2}\right)^{2}\left[w^{\top}0\cdots 0\right]^{\top},$ (7)
where $\delta^{\mathrm{Layer}}(\phi,\varsigma)$ is the TD update vector that adopts the LayerNorm critic—$Q_{\phi}^{\text{Layer }}(x)=w^{\top}$ LayerNorm [${}^{k}[\sigma\circ Mx]$.
D.3 Twin Delayed Deep Deterministic Policy Gradient (TD3)
Twin delayed deep deterministic policy gradient (TD3) (Fujimoto et al., 2018) learns a deterministic policy $\pi_{\bm{\theta}}(\bm{s})$ and two Q-functions $Q_{\bm{\phi}_{1}},Q_{\bm{\phi}_{2}}$ through clipped double Q-learning:
$y_{t}=r_{t}+\gamma\min_{i=1,2}Q_{\bm{\phi}_{i}^{-}}(\bm{s}_{t+1},\pi_{\bm{\theta}^{-}}(\bm{s}_{t+1})+\epsilon),$ (8)
where $\epsilon\sim\text{clip}(\mathcal{N}(0,\sigma),-c,c)$, with Q-function losses
$L_{Q}(\bm{\phi}_{i})=\mathbb{E}_{(\bm{s}_{t},\bm{a}_{t})\sim\mathcal{D}}[(Q_{\bm{\phi}_{i}}(\bm{s}_{t},\bm{a}_{t})-y_{t})^{2}],$ (9)
and policy loss
$L_{\pi}(\bm{\theta})=-\mathbb{E}_{\bm{s}_{t}\sim\mathcal{D}}[Q_{\bm{\phi}_{1}}(\bm{s}_{t},\pi_{\bm{\theta}}(\bm{s}_{t}))],$ (10)
employing delayed policy updates and target network stabilization.
Appendix E Benchmark Selection
Figure 2: Screenshots of well-established ALE, Procgen, DMC, and Craftax (left to right).
(a) Continual Procgen with cyclic task streams
(b) Continual Procgen with level shift
Figure 3: Maintaining plasticity in continual Procgen.
(a) Continual DMC with cyclic task streams
(b) Continual DMC with varying dynamic
Figure 4: Maintaining plasticity in continual DMC.
Benchmark Task Observation Space Action Space
ALE DemonAttack Box(0, 255, (210, 160, 3)) Discrete(6)
Procgen StarPilot Box(0, 255, (3, 64, 64)) Discrete(15)
DMC Dog Box(-inf, inf, (223,), float32) Box(-1.0, 1.0, (38,), float32)
Craftax Symbolic Box(0.0, 1.0, (8268,), float32) Discrete(43)
Table 1: Details of the challenging environments.
E.1 Standard
In the standard setting, as shown in Figure 2, we use four well-established and challenging benchmarks, i.e., Arcade Learning Environment (ALE) (Bellemare et al., 2013), Procgen (Cobbe et al., 2020), and DeepMind Control Suite (DMC) (Tassa et al., 2018).
E.2 Continual
E.2.1 Variant of Procgen
For Procgen, we evaluate the adaptability of various methods using two variants: one involves increasing the level with a fixed level offset (see Figure 3(a)), while the other involves constructing a task sequence, where each sequence is a random combination of tasks (see Figure 3(b)). It is important to note that we design multiple rounds for the effective evaluation of plasticity, with the task sequence in each round being randomized.
E.2.2 Variant of DMC
To validate the effectiveness of various plasticity-based methods, we use the Dog as the base agent and build a task stream including five different tasks, i.e., Stand, Walk, Trot, Run, Fetch (see Figure 4(a)). Akin to continual Procgen (task-shift), each round is assigned a random task stream. Moreover, to build a non-stationary DMC environment, as visualized in Figure 4(b), we change the dynamics of the Quadruped agent. Specifically, we varied the coefficient of friction by sampling it from a log-uniform distribution over the range [0.02, 2.00], applying a logarithmic scale with base 10.
E.3 Open-ended
To investigate the plasticity in an open-ended environment, we adopt Craftax (Matthews et al., 2024) as the test bunch, which evaluates the exploration and long-horizon capability of the agent.
Appendix F Plasticity-Related Evaluation Metrics
F.1 Ratio of Dormant Units
@@
F.2 Fraction of Active Units
@@ F.4 Effective Rank
F.5 Weight Magnitude
The weight magnitude is used to monitor the overall scale of the model parameters, which can provide insights into the model’s capacity usage and its level of plasticity (Lyle et al., 2024b). A high weight magnitude may suggest potential overfitting or reduced adaptability due to saturation effects, while a low magnitude might indicate undertraining or high plasticity.
Formally, denote by $f_{\Theta}$ the neural network with a set of parameters $\Theta$, the weight magnitude (WM) is defined as
$\text{WM}=\sqrt{\frac{\sum_{\theta\in\Theta}\theta^{2}}{|\Theta|}},$ (15)
where $|\Theta|$ is the cardinality.
F.6 Weight Difference
@@
Let $\boldsymbol{\theta}^{(t)}=\{\theta_{1}^{(t)},\dots,\theta_{P}^{(t)}\}$ denote the parameters of the model after learning task $t$, and $\boldsymbol{\theta}^{(t-1)}$ be the parameters before learning the task. The L2 norm difference between these two states is defined as (Juliani and Ash, 2024):
$\text{WD}=\sqrt{\frac{1}{P}\sum_{p=1}^{P}\left(\theta_{p}^{(t)}-\theta_{p}^{(t-1)}\right)^{2}}.$ (16)
This metric captures the average magnitude of change in parameters and helps assess whether the model underwent significant structural modification between training phases.
F.7 Feature Norm
The feature norm evaluates the average L2 norm of the feature vectors produced by the model, providing a quantitative measure of the feature magnitude (Kumar et al., 2022). This metric is particularly useful in continual reinforcement learning for detecting representational collapse or overly sparse activations, which may indicate reduced expressiveness or excessive regularization (Lee et al., 2024).
Given a feature matrix $\mathbf{F}\in\mathbb{R}^{n\times m}$, where $n$ is the batch size (number of samples) and $m$ is the dimensionality of each feature vector, the feature norm (FN) is computed as:
$\text{FN}(\mathbf{F})=\frac{1}{n}\sum_{i=1}^{n}\left\|\mathbf{F}_{i,:}\right\|_{2},$ (17)
where $\mathbf{F}_{i,:}$ denotes the $i$-th row of $\mathbf{F}$, corresponding to the feature representation of the $i$-th input.
F.8 Feature Variance
The feature variance quantifies the variability of feature activations across input samples. In continual reinforcement learning, this metric helps assess whether the model maintains diverse and informative representations over time. A collapse in feature variance may indicate overfitting, representational drift, or forgetting.
Let $\mathbf{F}\in\mathbb{R}^{n\times m}$ denote the feature matrix, where $n$ is the number of input samples and $m$ is the number of features. The feature variance (FV) is defined as the average variance across all feature dimensions (Lee et al., 2024):
$\text{FV}(\mathbf{F})=\frac{1}{m}\sum_{j=1}^{m}\operatorname{Var}(\mathbf{F}_{:,j}),$ (18)
where $\mathbf{F}_{:,j}$ denotes the $j$-th column of $\mathbf{F}$, corresponding to the $j$-th feature across the batch.
F.9 Gradient Norm
The gradient norm (GN) quantifies the magnitude of gradients during optimization, reflecting the learning signal’s strength (Abbas et al., 2023). In continual reinforcement learning, tracking the gradient norm is essential for diagnosing issues such as vanishing gradients, overfitting, or under-training. Moreover, it provides insights into how much the model updates in response to new data or tasks.
@@ F.9 Gradient Norm
F.10 Policy Entropy
Policy entropy quantifies the uncertainty or stochasticity of an agent’s policy (Haarnoja et al., 2018). In reinforcement learning, higher entropy encourages exploration by promoting diverse actions, while lower entropy indicates more deterministic (exploitative) behavior. In continual reinforcement learning, monitoring policy entropy helps analyze whether the agent becomes overconfident (risking forgetting) or maintains sufficient exploration across tasks.
Let $\pi(\bm{a}|\bm{s})$ denote the probability of selecting action $\bm{a}$ under state $\bm{s}$ according to the policy $\pi$. The entropy of the policy at a given state is defined as:
$H(\pi(\cdot|\bm{s}))=-\sum_{\bm{a}\in\mathcal{A}}\pi(\bm{a}|\bm{s})\log\pi(\bm{a}|\bm{s}),$ (20)
where $\mathcal{A}$ is the set of possible actions. Given a batch of states $\{\bm{s}_{1},\dots,\bm{s}_{n}\}$, the average policy entropy (PE) is:
$\text{PE}=\frac{1}{n}\sum_{i=1}^{n}H(\pi(\cdot|\bm{s}_{i})).$ (21)

### v2

@@ Abstract
Developing lifelong learning agents is crucial for artificial general intelligence (AGI). However, deep reinforcement learning (RL) systems often suffer from plasticity loss, where neural networks gradually lose their ability to adapt during training. Despite its significance, this field lacks unified benchmarks and evaluation protocols. We introduce Plasticine, the first open-source framework for benchmarking plasticity optimization in deep RL. Plasticine provides single-file implementations of over 13 mitigation methods, 6 evaluation metrics, and learning scenarios with increasing non-stationarity levels from standard to continually varying environments. This framework enables researchers to systematically quantify plasticity loss, evaluate mitigation strategies, and analyze plasticity dynamics across different contexts. Our documentation, examples, and source code are available at https://github.com/RLE-Foundation/Plasticine.
Keywords: Deep RL, Plasticity, Lifelong Learning, Benchmark, PyTorch
@@ 1 Introduction
Developing agents that never stop learning and evolving in dynamic real-world environments represents a critical frontier for advancing artificial intelligence and a fundamental prerequisite for continuously improving systems (Abel et al., 2023). Reinforcement learning (RL), with its inherent properties of learning from experience, continual online improvement, and iterative strategic exploration, emerges as the ideal framework for pursuing this objective (Silver and Sutton, 2025). Despite its promise, recent research has increasingly revealed a critical limitation: plasticity loss, a phenomenon in which neural networks in RL agents gradually lose their ability to adapt and incorporate new information as training progresses (Dohare et al., 2024; Klein et al., 2024), thus significantly impeding the development of truly lifelong learning agents (Lyle and Pascanu, 2024).
@@
Despite the clear importance of mitigating plasticity loss in deep RL, this field lacks unified methodological frameworks and evaluation protocols. This lack of established benchmarks hampers fair comparisons, methodological consistency, and precise problem formulation. Numerous fundamental questions remain inadequately addressed in current research: $\bullet$ How can we reliably measure and quantify plasticity loss through empirical means? $\bullet$ To what degree do existing mitigation approaches effectively resolve this issue? $\bullet$ In what ways does plasticity loss vary across different task domains and learning scenarios? To address these gaps, we introduce Plasticine, the first open-source benchmark for plasticity research in deep RL. The following section presents the framework architecture. Detailed method descriptions, experimental configurations, and results appear in the Appendices.
@@ 2 Architecture of Plasticine
(a) Architecture of the Plasticine framework. Vis. is short for Visual.
(b) Code implementation of the Plasticine-enhanced agent.
(c) Invocation of the Plasticine methods in the CleanRL workflow.
Figure 1: Overview of the Plasticine framework. Plasticine strikes a balance between “single-file” and “modularized” implementations, facilitating related research and redevelopment while effectively controlling the project size.
Figure 1 illustrates the overall architecture of the Plasticine framework, which contains three core parts: Methods, Metrics, and Environments. To maximize the convenience for researchers, we adopt the “single-file” implementation style that has been widely recognized by the RL community (Huang et al., 2022; Tarasov et al., 2023). However, as most plasticity enhancement methods are relatively small techniques that intervene in the agent’s network, adopting a purely single-file implementation style would lead to excessive project volume and management difficulties. Therefore, we strike a balance between single-file and modular design. As shown in Figure 1(a) and 1(b), all plasticity methods are integrated into the agent implementation, which is then combined with CleanRL’s single-file RL algorithm (Huang et al., 2022). This approach allows researchers to focus on plasticity methods and facilitates redevelopment, while effectively controlling the project size. The following sections detail the design insights.
@@ 2.1 Plasticity Loss Mitigation Methods
    Reset-based Intervention: Re-initializing portions of the network represents the most direct approach to regain plasticity (Nikishin et al., 2022), whether through manually selecting specific layers or automatically identifying neurons experiencing severe performance plateaus (Sokar et al., 2023; Galashov et al., 2024).
@@
    Optimizer: Specialized optimization approaches like Fast TRAC (Muppidi et al., 2024) and Kron (Castanyer et al., 2025) help navigate complex loss landscapes during non-stationary training, thereby maintaining plasticity.
@@ 2.2 Comprehensive Evaluation Metrics
2.3 Progressive Evaluation Scenarios
@@
2.3.1 Standard Online RL
Standard online RL naturally exhibits non-stationarity as policies evolve and through bootstrapping processes. In these environments, agents must maintain plasticity throughout training to achieve optimal performance instead of being stuck at suboptimal levels. This setting offers a foundation for measuring plasticity during natural distribution shifts in policy optimization without introducing explicit task changes. Our baseline evaluation includes the Arcade Learning Environment (ALE; Bellemare et al., 2013) as the benchmark, which has been increasingly employed to study plasticity loss in deep RL (Kumar et al., 2021, 2022; Sokar et al., 2023; Castanyer et al., 2025).
2.3.2 Continual RL
Continual RL scenarios create more challenging plasticity demands by introducing explicit distribution shifts beyond those naturally occurring in standard RL. Following the practice in (Muppidi et al., 2024; Tang et al., 2025), we consider two scenarios: one in which visual features change within the same task, and another that requires adaptation across entirely different tasks. These environments test an agent’s ability to maintain plasticity across both sudden task switches and gradual dynamics changes, providing a more challenging testbed than standard online RL. With clearly defined non-stationarity boundaries, we can directly observe how models behave before and after distribution shifts, gaining valuable insights into plasticity degradation patterns across different task distributions.
3 Conclusion
Maintaining plasticity is essential for building agents that learn continually throughout their lifetimes, a central goal as reinforcement learning enters the era of experience. We present Plasticine, an open-source framework designed to accelerate research in this direction. Plasticine offers three key advantages: (1) ease of use through single-file implementations built on familiar RL workflows; (2) broad coverage spanning on-policy and off-policy algorithms, state-based and vision-based domains, and evaluation scenarios from standard online RL to continual settings; and (3) comprehensive analysis tools for diagnosing plasticity dynamics across methods and environments. We will continue to expand Plasticine with new methods and environments, and hope it serves as a useful foundation for future research in this rapidly evolving field.
Appendix A Supplementary Material
@@
  • Appendix D provides formally mathematical definitions of the plasticity-related metrics;
  • Appendix E introduces the implementation details of the training environments for the two learning scenarios (standard and continual RL);
  • Appendix F introduces the included backbone RL algorithms;
@@ Appendix B Project Structure
Plasticine/
    plasticine/ (Main algorithm implementation)
    ppo_continual_procgen_base.py (torch.nn.Module-based agent integrated with diverse plasticity loss-mitigation techniques)
    ppo_continual_procgen_plasticine.py (Combination of the Plasticine agent and the CleanRL workflow)
    utils.py (Auxiliary functions)
    plasticine_envs/ (Environments for algorithm evaluation)
    procgen_wrappers.py (Continual Procgen benchmark)
    dmc_wrappers.py (Continual DMC benchmark)
    plasticine_metrics/ (Plasticity-related evaluation metrics)
    metrics.py
    requirements/ (Dependencies for environments)
    requirements-ale.txt
    requirements-prcogen.txt
    scripts/ (Training scripts)
    experimental/ (Experimental code files)
    assets/ (Miscellaneous files)
@@ C.5.1 TRAC
C.5.2 KRON
KRON (Castanyer et al., 2025) is a Kronecker-factored optimizer that addresses gradient pathologies in deep RL by providing curvature-aware updates. Unlike first-order optimizers like Adam that rely on diagonal preconditioning, Kron approximates the Fisher information matrix using Kronecker products, enabling more precise adaptation to the loss landscape’s curvature. This is particularly crucial in non-stationary RL settings, where policy-dependent data distributions and bootstrapped targets lead to unstable optimization trajectories.
Appendix D Plasticity-Related Evaluation Metrics
D.1 Ratio of Dormant Units
@@
D.2 Fraction of Active Units
@@ D.4 Effective Rank
D.5 Weight Difference
@@
D.6 Gradient Norm
The gradient norm (GN) quantifies the magnitude of gradients during optimization and reflects the strength of the learning signal (Abbas et al., 2023). In continual reinforcement learning, tracking the gradient norm is essential for diagnosing issues such as vanishing gradients, overfitting, or under-training. Moreover, it provides insights into how much the model updates in response to new data or tasks.
@@ D.6 Gradient Norm
Appendix E Benchmark Selection
E.1 ALE
We first introduce the Arcade Learning Environment (ALE) as the benchmark, which is a collection of arcade game environments that require the agent to learn motor control directly from image-based observations. This benchmark has been used to examine the loss of plasticity in a series of continual RL studies (Kumar et al., 2021, 2022; Sokar et al., 2023; Castanyer et al., 2025). As shown in Figure 2, the ALE environments often integrate distinct levels or stages. The transitions between these levels introduce significant changes in visual appearance, dynamics, or task objectives. Such level-wise shifts further exacerbate the non-stationarity of the learning problem, thereby leading to a significant challenge in terms of plasticity loss. In our experiments, we stack four consecutive frames to form an input state with the data shape of $(84,84,4)$ for all environments.
Figure 2: The ALE benchmark with visual observations and discrete action space.
E.2 Continual Procgen
Figure 3: Continual Procgen benchmark with visual observations and discrete action space.
Furthermore, we introduce the Procgen benchmark, which comprises 16 procedurally generated environments. Procgen is developed similarly to the ALE benchmark, in which agents must learn motor control directly from image-based observations. However, Procgen involves much higher dynamics that require the agent to continually learn and adapt to the environment. Following the practice in (Muppidi et al., 2024) and (Tang et al., 2025), we implement a variant of Procgen to support the comprehensive evaluation of the implemented methods. As shown in Figure 3, distribution shifts are introduced by sampling a new procedurally generated level of the current game every 2 million time steps, and each level is considered a distinct task. In our experiments, all environments use a discrete fifteen-dimensional action space and generate $(64,64,3)$ RGB observations, with the easy mode employed.
E.3 Continual DMC
Figure 4: Continual DMC benchmark with proprioceptive observations and continuous action space.
In addition to the discrete-action settings above, we further implement a variant of the DeepMind Control Suite (DMC) benchmark (Tassa et al., 2018) following (Tang et al., 2025), which contains a series of continuous control tasks. Notably, each environment is a continual RL scenario that chains the corresponding individual tasks in a sequence. As illustrated in Figure 4, the Continual Dog environment chains four individual tasks: Dog Stand, Dog Walk, Dog Run, and Dog Trot. For each individual task, we train it for 1M environment steps.
Appendix F Backbone RL Algorithms
In this paper, we select two representative RL baselines as the backbone algorithms, which span policy-based, value-based, on-policy, and off-policy approaches.
F.1 Proximal Policy Optimization
Proximal policy optimization (PPO) (Schulman et al., 2017) is a popular on-policy algorithm for learning a continuous or discrete control policy $\pi_{\bm{\theta}}(\bm{a}|\bm{s})$. PPO employs policy gradients using action-advantages: $A_{t}=A^{\pi}(\bm{a}_{t},\bm{s}_{t})=Q^{\pi}(\bm{a}_{t},\bm{s}_{t})-V^{\pi}(\bm{s}_{t})$, and minimizes a clipped-ratio loss over mini-batches of recent experience (collected under $\pi_{\bm{\theta}_{\text{old}}}$):
[…]
F.2 Categorical DQN
Categorical DQN (C51) (Bellemare et al., 2017) is a distributional RL algorithm that models the return distribution rather than its expectation. Instead of directly learning the action-value function $Q^{\pi}_{r}(\bm{s},\bm{a})$, C51 learns a categorical distribution of the random return $Z(\bm{s},\bm{a})$ on a fixed support $\{z_{i}\}_{i=1}^{N}$, where
$z_{i}=V_{\min}+(i-1)\Delta z,\Delta z=\frac{V_{\max}-V_{\min}}{N-1},\ \ i=1,\dots,N.$ (12)
The probability of atom $z_{i}$ is $p_{\bm{\theta}}(z_{i}\mid\bm{s},\bm{a})=\frac{\exp(f_{i}(\bm{s},\bm{a};\bm{\theta}))}{\sum_{j}\exp(f_{j}(\bm{s},\bm{a};\bm{\theta}))}$. Here, $f_{i}(\bm{s},\bm{a};\bm{\theta})$ denotes the logit for atom $z_{i}$, $\bm{\theta}$ is a parametric model. Given a sampled transition $(\bm{s},\bm{a},r,\bm{s^{\prime}})$, C51 forms the distributional Bellman target using the target network $\bm{\theta}^{-}$, then projects it back onto the fixed support via $\Phi$. We minimize the Kullback–Leibler (KL) divergence between the projected target and the predicted distribution:
$L(\bm{\theta})=D_{\mathrm{KL}}\!\left(\Phi\hat{\mathcal{T}}Z_{\bm{\theta}^{-}}(\bm{s},\bm{a})\ \|\ Z_{\bm{\theta}}(\bm{s},\bm{a})\right)=-\sum_{i=1}^{N}m_{i}\log p_{\bm{\theta}}(z_{i}\mid\bm{s},\bm{a}),$ (13)
where $m=\Phi\hat{\mathcal{T}}Z_{\bm{\theta}^{-}}(\bm{s},\bm{a})$, and $\hat{\mathcal{T}}$ denotes the distributional Bellman optimality operator.